A company issued 6-year, 8% bonds with a par value of $850,000. The market rate when the bonds were issued was 7.5%. The company received $858,500 cash for the bonds. Using the straight-line method, the amount of recorded interest expense for the first semiannual interest period is:
Answer #1

Amortization of premium for a semiannual period

= (858,500-850,000)/12 = 708.33

Interest expense = 850,000 * 8% * 6/12 = 34,000

= 34,000 - 708.33

= 33,292

Comment if you face any issues